OpenStudy (anonymous):

Use completing the square to find the maximum height of the projectile: -16t^2+64t+80

OpenStudy (campbell_st):

ok... so you are being asked to find the vertex of the parabola, by putting it into vertex form. so \[y = -(16t^2 - 64t + ?) + 80\] so \[16t^2 = (4t)^2 \] so you need to find the value of the the last term remember (a - b)^2 = a^2 - 2ab + b^2 you know a = 4t find b

OpenStudy (anonymous):

how would I find b? plug in 4t into A?